# Spurious reconstruction from brain activity

### v3

@@ Results
We first confirmed the reproduction of the findings of the text-guided reconstruction methods, particularly those of the Brain-Diffuser study [45] when applied to the NSD dataset (Fig 2A). The reconstructed images generated by the Brain-Diffuser method effectively captured most of the layout and semantics of the test images. Although the reconstructed images produced by the StableDiffusionReconstruction method [44] showed slightly degraded performance compared to the original paper, they still successfully captured the semantics of the test images. Notably, the iCNN method [29] also performed well on the NSD dataset, with reconstructed images capturing the dominant structures of the objects within the images, consistent with the findings reported in the original study.
@@
Fig 2: Comparison of image reconstruction results across datasets and methods. (A) Reconstruction results for the Natural Scene Dataset (NSD). The first row shows the original test images, followed by reconstructions using StableDiffusionReconstruction, Brain-Diffuser, and iCNN methods. (B) Reconstruction results for the Deeprecon dataset, presented in the same format as (A).
@@
To further investigate the generalizability of the text-guided reconstruction methods, we attempted to replicate their performance using a different dataset, Deeprecon, which was originally collected for the study by Shen et al. [29]. The Deeprecon dataset was explicitly designed to avoid overlap between training and test sets, making it an ideal benchmark for evaluating the zero-shot prediction capabilities of reconstruction methods. However, the original Deeprecon dataset lacked the text annotations required by the text-guided reconstruction methods. To facilitate a fair comparison, we collected five text annotations for each training stimulus in the Deeprecon dataset through crowd-sourcing and used them to generate CLIP text features.
@@
Despite including these text annotations, the text-guided reconstruction methods failed to achieve the same level of performance on the Deeprecon dataset as they did on the NSD dataset. The reconstructed images produced by the text-guided methods exhibited realistic appearances but suffered from largely degraded quality compared to their performance on the NSD dataset (Fig 2B). Notably, the text-guided reconstruction methods generated realistic reconstructions even for simple geometric shapes in the Deeprecon dataset, which deviated significantly from the original stimuli. In contrast, the iCNN method consistently provided faithful reconstructions for both the NSD and Deeprecon datasets, despite its simplicity compared to the text-guided methods. These results suggest that the text-guided methods may not generalize well across different datasets. The ability to generate realistic but inaccurate reconstructions, especially for simple shapes, indicates that these methods might rely more on training stimuli than on actual brain activity information.
@@
Upon further investigation, we noted a questionable post-hoc image selection procedure. In the StableDiffusionReconstruction study [44], they presented the reconstruction results by the following procedure: “We generated five images for each test image and selected the generated images with highest PSM” (PSM means perceptual similarity metrics, and it was calculated from early/middle/late layers of several image recognition deep neural networks [DNNs]), as illustrated in Fig 3A. This selection might lead readers or peer reviewers, particularly those not specialized in the brain decoding field, to overestimate the effectiveness of the methods and potentially lead to a distorted understanding of the actual reconstruction performance. Note that the Brain-Diffuser study [45] did not execute such procedures, and in their subsequent report [53], they updated the image presentation procedure more fairly as: “we generated five images with different stochastic noise and selected three images randomly.”
@@
Fig 3: Analysis of post-hoc selection in the StableDiffusionReconstruction method [44]. (A) Selection procedure. Five images were generated, and the one most closely resembling the test image was selected based on perceptual similarity metrics (PSM). (B) Images were generated from the CLIP text features translated from original brain activity and the VAE (vision) features from shuffled brain activity. Examples selected from five generations are shown with the test images. (C) Images were generated from the CLIP text and the VAE (vision) features, both translated from shuffled brain activity. Examples of selected images are shown with the text images. Examples are shown as in (B).
@@
To examine the impact of this post-hoc selection procedure, we conducted an experiment using random brain data. Instead of feeding the test brain data into trained translators, we shuffled the activities within each voxel of the NSD test set independently to create random brain data. Specifically, the brain data of the NSD test set is in a matrix shape, with rows representing stimulus samples and columns representing voxels. To generate the random brain data, we selected one column (voxel) of the matrix and randomly shuffled its values. This process was repeated for all voxels independently. Surprisingly, when the random brain data were input into the VAE feature translator, which contributes to producing initial images, plausible images were obtained by generating five images and selecting the best one (Fig 3B). Even more strikingly, when the random brain data were input into both VAE and CLIP text feature translators, we still obtained convincing results by simply generating images five times and conducting the selection mentioned above (Fig 3C). These observations are inexplicable because the artificially created brain data should completely lack any information related to the original visual stimuli.
@@
These observations raise perplexing questions about the performance and generalizability of recent text-guided reconstruction methods. The performance significantly deteriorated when we changed from the NSD to Deeprecon dataset. This highlights the need to better understand why these methods succeed with the NSD. Moreover, the ability to obtain plausible reconstructions from random brain data by merely generating multiple images and selecting the best ones suggests that there may be fundamental issues with both the evaluation dataset and the components of the reconstruction methods themselves. In the following sections, we will thoroughly investigate the potential problems associated with the NSD dataset and each component of the text-guided reconstruction pipeline.
Lack of diversity in the stimulus set
@@
First, we examined the characteristics and limitations of the NSD dataset itself. To characterize the diversity of stimuli in the datasets and their latent representations, we focused on the CLIP text features, which are used as latent features in text-guided reconstruction methods. We employed uniform manifold approximation and projection (UMAP) [54] to visualize the CLIP text features of the NSD stimuli (see Materials and Methods “UMAP visualization”). The visualization revealed approximately $40$ distinct clusters, with considerable overlap between the training and test sets (Fig 4A). Interestingly, we were able to describe the stimulus images in each cluster using a single semantic word, such as airplane, giraffe, or tennis. Despite the NSD containing around $30,000$ brain samples per subject, the diversity of the presented stimuli was quite limited to just around $40$ semantic concepts. Note that we performed UMAP visualization using the parameters recommended in the official guide for clustering, but a similar cluster structure was also observed with the default parameters (S1 Fig). In contrast, the Deeprecon dataset, which was specifically designed to differentiate object categories between training and test data, exhibited less overlap between the two sets (S2 Fig).
@@
Fig 4: Dataset diversity and similarity between training and test stimuli. (A) UMAP visualization of CLIP text features in the NSD. The center figure shows the scatter plot of the UMAP embedding of CLIP text features. The gray points represent training samples, while the orange points represent test samples. The surrounding images were randomly selected from each cluster. (B, C) Similarity between training and test images. For each example test image (the one on the left of each row), the five training images with the highest similarity were selected using the DreamSim metric and displayed. This analysis was performed for the NSD (B) and Deeprecon (C) datasets, using the same procedure.
@@
To further investigate the similarity between the training and test stimuli, we analyzed using a state-of-the-art perceptual similarity metric called DreamSim [55]. For each test image in the NSD, we used DreamSim to identify the most similar training images based on their perceptual similarity scores. We found that the training images extracted using DreamSim were highly similar to the test images in the NSD, not only in terms of semantic concepts but also in terms of overall layout and visual composition (Fig 4B). These findings suggest that the distribution of NSD test images is heavily biased towards that of the training images, with a significant overlap in the visual and semantic features present in both sets. Such a strong bias raises concerns about the actual reconstruction performance of methods evaluated on this dataset. The impressive results may be largely attributable to the methods’ ability to memorize and reproduce the characteristics of the training stimuli rather than generalizing them to novel stimuli.
In contrast, when we applied the same DreamSim-based analysis to the Deeprecon dataset, we found that the extracted training images were substantially different from the test images (Fig 4C). This observation supports the careful design of the train-test split in the Deeprecon dataset, ensuring that the test images contain novel visual and semantic features that are not well-represented in the training set. The distinct differences between the NSD and Deeprecon datasets in terms of stimulus similarity highlight the importance of carefully designing evaluation benchmarks that can effectively assess the generalization capabilities of visual image reconstruction methods.
Failed zero-shot prediction in the feature space
Given the significant semantic and structural similarities between the NSD test images and the training images, it is crucial to assess whether the predictions derived from brain activity truly demonstrate generalization beyond the training set (i.e., zero-shot prediction). To evaluate the zero-shot prediction capability of the translators used in the reconstruction methods, we conducted an $(N+1)$-way identification analysis, where $N$ represents all of the training samples ($8,859$ samples for subject 1 in the NSD) and $1$ refers to the test sample (Fig 5A). Specifically, we calculated the similarity (correlation) between the predicted (translated) features and the true features of corresponding test samples, as well as the similarities between the predicted features and the true features corresponding to each of the N training samples. We then examined whether the predicted features correctly identified (showed the highest similarity to) the test true features. We calculated the percentage of predicted features that successfully identified the corresponding test sample in the NSD dataset.
@@
Fig 5: Evaluation of zero-shot sample identification performance. (A) Schematic diagram of novel sample identification. Similarity is calculated between predicted (translated) features and test features, as well as between predicted features and each of the training features. Identification accuracy is determined by how often the predicted features show the highest similarity to the true test features compared to all training features. (B) Zero-shot identification performance using predicted CLIP features (text and vision) from the Brain-Diffuser method, evaluated on the NSD dataset. (C) Zero-shot identification performance using predicted features from several intermediate layers of the VGG19 model employed in the iCNN method, also evaluated on the NSD dataset. For both (B) and (C), the identification task involved $8,859$ training samples plus one test sample, resulting in an $8,860$-way identification. The chance performance level is thus $1/8860$.
The performance of the CLIP features was nearly 00% (Fig 5B). The CLIP features predicted from brain activity were unable to accurately identify novel test samples, even within the distribution covered by the training set. This extremely low performance suggests that the CLIP feature translator fails to achieve zero-shot prediction. This finding raises questions about the effectiveness of CLIP features for true zero-shot prediction tasks in brain decoding. In contrast, the VGG19 features [56] employed in the iCNN method demonstrated moderate identification performance at the latent features on the intermediate layers (Fig 5C). This finding suggests that the VGG19 features possess a certain degree of generalizability, enabling them to identify test samples not encountered during training.
@@
Furthermore, we investigated whether the CLIP feature translator enables the prediction of novel semantic concepts not in the training. We redesigned the dataset split to ensure no semantic concepts were shared between them as in the previous zero-shot prediction study [17, 18]. We first applied $k$-means clustering to the UMAP embedding space of the NSD’s CLIP text features (S3 FigA). We set the number of clusters as $40$ based on visual inspection of the UMAP results. Based on these clustering results, we performed a hold-out analysis: when predicting samples within a cluster (e.g., the ski cluster), we excluded samples from that cluster in the training set (Fig 6A right; hold-out split condition). As a control, we also prepared a naive data-split condition where the training sample size is the same as in the hold-out split condition but allows overlapping semantic concepts (Fig 6A left; naive split condition). When we visualized the properties of predicted features in hold-out analysis by transforming them into the previous UMAP embedding space (Fig 4A), we observed that the predicted features tend to diverge largely from their original clusters and move into other clusters (Fig 6B).
@@
Fig 6: Cluster hold-out analysis. (A) Hold-out procedure. When predicting a test sample from a specific cluster, all training samples from that cluster are excluded. The naive split condition serves as a control, using the same number of training samples but allowing cluster overlap between training and test sets. (B) Examples of predictions. Under the hold-out split condition, gray and dark orange points represent training and test samples, respectively. Light orange points indicate translated (predicted) latent features. Black lines connect true features to their corresponding predictions. (C) Cluster identification procedure. The predicted features are compared with the average features of each cluster, and the most similar cluster is selected. (D) Cluster identification results. Surrounding panels display CLIP text and vision feature performance in the hold-out and naive split conditions of the four representative semantic clusters. The chance level is $1/40$. (E) Pairwise sample identification procedure. Predicted features are compared with the true test sample and another randomly chosen test sample. A sample more similar to the prediction is selected. (F) Pairwise sample identification results. The accuracy means the average winning rate of the true samples against all of the other test samples. Results are presented similarly to cluster identification, with a chance level of $1/2$.
@@
To quantitatively assess the performance of the feature translator, we employed two identification metrics. The first metric is cluster identification accuracy. Cluster identification accuracy focuses on evaluating the translator’s ability to generate features that correctly identify the semantic cluster to which a test sample belongs (Fig 6C). In this analysis, we compare the predicted features of each test sample to the average features of the training samples within each semantic cluster. The accuracy is then calculated as the percentage of predicted features that successfully identify the original semantic cluster of their corresponding test samples in the latent feature space.
@@
The cluster identification accuracy in the hold-out split condition exhibited a substantial drop compared to the naive split condition across all semantic clusters (Fig 6D). Notably, the cluster identification accuracy was frequently 00% in the hold-out split condition (see also S3 Fig.B for all cluster results). These results expose a severe limitation of the translator when dealing with novel semantic concepts absent from the training set. This suggests that the CLIP feature translator primarily functions as a “classifier;” its prediction (translation) heavily relies on predefined semantic features used in a training phase rather than generalizing to new concepts.
@@
The second metric is pairwise sample identification accuracy. Pairwise sample identification accuracy is a commonly used metric in the evaluation of feature prediction and reconstruction performance [39, 29, 34, 36, 40, 41, 44, 45, 46, 57, 58]. This analysis assesses the translator’s ability to predict target features that are more similar to the actual test features than to the features of other samples in the test set (Fig 6E). For each test sample, we compare its predicted features with two sets of features: the actual test features and the features of another randomly selected sample from the test set. The accuracy is then calculated as the average winning rate of the predicted features against all other test samples, indicating the proportion of cases where the predicted features are more similar to the actual test features than those of other samples.
Intriguingly, despite the feature translator’s complete failure to identify semantic concepts in the hold-out split condition, the pairwise sample identification accuracy often surpassed the chance level across the semantic clusters (Fig 6F). The insensitivity of the pairwise sample identification accuracy to the translator’s inability to capture semantic concepts can be explained by its focus on relative similarity. Specifically, the similarity compares the predicted features to the actual test features, without taking into account the broader general context. This can lead to an inflated pairwise sample identification accuracy, potentially misrepresenting the actual reconstruction performance and the translator’s ability to generalize to novel clusters. This observation raises concerns about the adequacy of using identification accuracy as the sole metric for evaluating reconstruction performance, a practice that has been widely adopted in many studies.
Failed recovery of a stimulus from its latent features

### v5

@@ 2.1.1 Observations: Failed replication and convincing reconstruction from random data
We first confirmed the reproduction of the findings of the original methods (Fig. 2A). The reconstructed images produced by the StableDiffusionReconstruction method (Takagi and Nishimoto, 2023a) showed slightly degraded performance compared to the original paper, but still successfully captured the semantics of the test images. The Brain-Diffuser method (Ozcelik and VanRullen, 2023) effectively captured most of the layout and semantics of the test images when applied to the NSD dataset. Similarly, MindEye2 (Scotti et al., 2024) generated reconstructions that preserved key visual elements of the original stimuli. Notably, despite originally being validated on a different dataset (Deeprecon), the iCNN method (Shen et al., 2019b) also performed well on the NSD dataset, capturing the dominant structures of the objects. This performance is consistent with the findings reported in the original study.
@@
Fig. 2: Comparison of image reconstruction results across datasets and methods. (A) Reconstruction results for the Natural Scene Dataset (NSD). The first row shows the original test images, followed by reconstructions using StableDiffusionReconstruction, Brain-Diffuser, MindEye2, and iCNN methods. (B) Reconstruction results for the Deeprecon dataset, presented in the same format as (A).
@@
To further investigate the generalizability of the text-guided reconstruction methods, we attempted to replicate their performance using a different dataset, Deeprecon, which was originally collected for the study by Shen et al. (2019b). The Deeprecon dataset was explicitly designed to avoid overlap between training and test sets, making it a suitable benchmark for evaluating the zero-shot prediction capabilities of reconstruction methods. However, the original Deeprecon dataset lacked the text annotations required by the text-guided reconstruction methods. To enable a fair comparison, we collected five text annotations for each training stimulus in the Deeprecon dataset through crowd-sourcing and used them to generate CLIP text features.
@@
Despite our careful replication efforts, including preparing text annotations, both text-guided reconstruction methods and MindEye2 failed to achieve the same level of performance on the Deeprecon dataset as they did on the NSD dataset (Fig. 2B). The reconstructed images produced by the text-guided methods exhibited realistic appearances but suffered from largely degraded quality compared to their performance on the NSD. Notably, the text-guided reconstruction methods generated realistic reconstructions even for simple geometric shapes in the Deeprecon dataset, which deviated strikingly from the original stimuli. Similarly, the reconstructions from MindEye2 did not resemble the test images but tended to exhibit object categories in the Deeprecon training set (frog, fighter aircraft, baby buggy, or baseball glove). Sample size matching between NSD and Deeprecon yielded similar reconstruction quality (Fig. A1), suggesting that sample size alone does not account for the poorer results on the Deeprecon. In contrast, the iCNN method consistently provided faithful reconstructions for both the NSD and Deeprecon datasets, despite its simplicity compared to these methods. These results suggest that the text-guided reconstruction methods and MindEye2 struggle to generalize across different datasets. The tendency to generate realistic yet inaccurate reconstructions, especially for simple shapes, indicates that these methods might rely more on learned training stimuli than on actual brain activity information.
@@
Upon further investigation, we noted a questionable post-hoc image selection procedure. In the StableDiffusionReconstruction study (Takagi and Nishimoto, 2023a), they presented the reconstruction results by the following procedure: “We generated five images for each test image and selected the generated images with highest PSM.” In their paper, PSM refers to perceptual similarity metric, which was calculated from early, middle, and late layers of several image recognition DNNs. This procedure is illustrated in Fig. 3A. This selection might lead readers or peer reviewers, particularly those not specialized in the brain decoding field, to overestimate the effectiveness of the methods and potentially lead to a distorted understanding of the actual reconstruction performance. Note that the Brain-Diffuser and MindEye2 studies did not execute such procedures, and in their subsequent report (Takagi and Nishimoto, 2023b), they updated the image presentation procedure more fairly as: “we generated five images with different stochastic noise and selected three images randomly.”
@@
Fig. 3: Analysis of post-hoc selection in the StableDiffusionReconstruction method (Takagi and Nishimoto, 2023a). (A) Selection procedure. Five images were generated, and the one most closely resembling the test image was selected based on perceptual similarity metrics (PSM). (B) Images were generated from the CLIP text features translated from original brain activity and the VAE (vision) features from shuffled brain activity. Examples selected from five generations are shown with the test images. (C) Images were generated from the CLIP text and the VAE (vision) features, both translated from shuffled brain activity. Examples of selected images are shown with the text images. Examples are shown in (B).
@@
To examine the impact of this post-hoc selection procedure, we conducted an experiment using random brain data. Instead of feeding the test brain data to trained translators, we shuffled the activities within each voxel of the NSD test set independently to create random brain data. Specifically, the brain data of the NSD test set is in a matrix shape, with rows representing stimulus samples and columns representing voxels. To generate the random brain data, we selected one column (voxel) of the matrix and randomly shuffled its values. This process was repeated for all voxels independently. Surprisingly, when the random brain data were input into the VAE feature translator, which contributes to producing initial images, plausible images were obtained by generating five images and selecting the best one (Fig. 3B). Even more strikingly, when the random brain data were input into both VAE and CLIP text feature translators, we still obtained convincing results by simply generating images five times and conducting the selection mentioned above (Fig. 3C). These observations are inexplicable because the artificially created brain data should completely lack any information related to the original visual stimuli.
@@
These observations raise perplexing questions about the performance and generalizability of recent reconstruction methods. Their performances largely deteriorated when we switched the evaluated dataset from the NSD to Deeprecon, highlighting the need to better understand why these methods succeed with the NSD. Moreover, the ability to obtain plausible reconstructions from random brain data by merely generating multiple images and selecting the best ones suggests that there may be fundamental issues with both the evaluation dataset and the components of the reconstruction methods themselves. In the following sections, we will thoroughly investigate the potential problems associated with the NSD dataset and each component of the text-guided reconstruction pipeline.
2.1.2 Lack of diversity in the stimulus set
@@
First, we examined the characteristics and limitations of the NSD dataset itself. To characterize the diversity of stimuli in the datasets and their latent representations, we focused on the CLIP text features, which are used as latent features in text-guided reconstruction methods. We employed uniform manifold approximation and projection (UMAP) (McInnes et al., 2018) to visualize the CLIP text features of the NSD stimuli (see Methods “UMAP visualization”). The visualization revealed approximately $40$ distinct clusters, with considerable overlap between the training and test sets (Fig. 4A). Interestingly, we were able to describe the stimulus images in each cluster using a single semantic label, such as airplane, giraffe, or tennis. Despite the NSD containing around $30,000$ brain samples per subject, the diversity of the presented stimuli was quite limited to just around $40$ semantic categories. Here, we performed UMAP visualization using the parameters recommended in the official guide for clustering. Even with the default UMAP parameters, a similar cluster structure was observed (Fig. A2). In contrast, the Deeprecon dataset, which was specifically designed to differentiate object categories between training and test data, exhibited less overlap between the two sets (Fig. A3; see Fig. A4 for the latent features used in MindEye2).
@@
Fig. 4: Dataset diversity and similarity between training and test stimuli. (A) UMAP visualization of CLIP text features of the NSD stimuli. The center figure shows the scatter plot of the UMAP embedding of CLIP text features. The gray points represent training samples, while the orange points represent test samples. The surrounding images were randomly selected from each cluster. (B, C) Similarity between training and test images. For each example test image (the one on the left of each row), the five training images with the highest similarity were selected using the DreamSim metric and displayed. This analysis was performed for the NSD (B) and Deeprecon (C) datasets, using the same procedure.
To further investigate the similarity between the training and test stimuli, we analyzed stimulus images using DreamSim, a state-of-the-art perceptual similarity metric (Fu et al., 2024). DreamSim was used to identify the most perceptually similar training images for each test image. We found that the training images identified by DreamSim metric were highly similar to the test images in the NSD, not only in semantic labels but also in overall layout and visual composition (Fig. 4B). In contrast, the same analysis on the Deeprecon dataset revealed that its training images were substantially different from the test images (Fig. 4C). To further assess whether the similarity between NSD training and test images is excessively high, we measured the similarity between the two sets. As a reference, we also measured the similarity between NSD training images and a large-scale independent dataset (CC3M; Sharma et al., 2018). This analysis revealed that the NSD test set contained stimuli that were highly similar to the training set, compared to those in the independent dataset (Fig. A5). In contrast, the Deeprecon test set, due to its carefully designed training–test split, exhibited a similar level of similarity to the independent dataset.
@@
These findings suggest that the distribution of NSD test images is heavily biased toward that of the training images, with a significant overlap in the visual and semantic features present in both sets. Such a strong bias raises concerns about the actual reconstruction performance of methods evaluated on this dataset. The convincing reconstruction results with the NSD may be largely attributed to the methods’ tendency to replicate specific characteristics observed in the training set, potentially at the expense of generalizing to novel stimuli. The distinct differences between the NSD and Deeprecon datasets in terms of stimulus similarity highlight the importance of carefully designing evaluation benchmarks to rigorously evaluate the generalization capabilities of visual image reconstruction studies.
2.1.3 Failed zero-shot prediction in the feature space
Given that many of the NSD test images closely resemble those in the training set, it is uncertain whether the translator’s predictions genuinely capture new, unseen stimuli (i.e., zero-shot generalization) or simply replicate features of training images that share similar semantics and visual structure. To probe this, we evaluated the zero-shot prediction capability of the CLIP feature translators by conducting an $(N+1)$-way identification analysis, where $N$ represents the entire training set ($8,859$ samples for subject 1 in the NSD) and $1$ is the target test sample (Fig. 5A). Concretely, we measured the correlation between each translator’s predicted features and the true features of (1) the correct test sample and (2) every training sample. We then asked whether the predicted features were most similar to the true test features, thereby correctly identifying them. Identification performance above chance suggests that the translator captures information specific to the test sample beyond the learned training patterns, supporting zero-shot prediction. Conversely, poor performance indicates that the translator does not predict the unique properties of the test sample.
@@ 2.1.3 Failed zero-shot prediction in the feature space
Fig. 5: Evaluation of zero-shot sample identification performance. (A) Schematic diagram of novel sample identification. Similarity is calculated between predicted (translated) features and test features, as well as between predicted features and each of the training features. Identification accuracy is determined by how often the predicted features show the highest similarity to the true test features compared to all training features. (B) Zero-shot identification performance using predicted CLIP features (text and vision) from the Brain-Diffuser method, evaluated on the NSD dataset. (C) Zero-shot identification performance using predicted features from several intermediate layers of the VGG19 model employed in the iCNN method, also evaluated on the NSD dataset. For both (B) and (C), the identification task involved $8,859$ training samples plus one test sample, resulting in an $8,860$-way identification. The chance performance level is thus $1/8860$.
The results show that the identification performance of CLIP feature translators was nearly 00% (Fig. 5B). This poor performance suggests that the CLIP feature translator captures only rough semantic categories encountered in the training set rather than fine-grained, instance-level details. This finding raises a question about the effectiveness of CLIP features for zero-shot prediction tasks in brain decoding.
By contrast, the VGG19 features (Simonyan and Zisserman, 2015), as used in the iCNN method, demonstrated moderate identification performance at the intermediate layers (Fig. 5C). This success could be attributed to the compositional representation of VGG19’s intermediate features. Unlike CLIP features, which are optimized for semantic alignment across vision and language, VGG19’s intermediate features, extracted through convolutional layers, contain primarily visual representations that retain sufficient local structure. This characteristic allows the translator to generalize to novel images by combining learned local spatial features rather than relying solely on semantic similarity. Consequently, VGG19’s compositional features may help achieve a certain degree of zero-shot prediction, distinguishing new images from closely resembling training examples.
@@
Furthermore, we investigated whether the CLIP feature translator enables the prediction of novel semantic clusters not in the training. We redesigned the dataset split to ensure no semantic clusters were shared between them as in previous zero-shot prediction studies (Mitchell et al., 2008; Brouwer and Heeger, 2009). We first applied $k$-means clustering to the UMAP embedding space of the NSD’s CLIP text features (Fig. A6A). We set the number of clusters as $40$ based on our visual inspection of the UMAP results. Based on these clustering results, we performed a hold-out analysis: when predicting samples within a cluster (e.g., ski cluster), we excluded samples of that cluster from the training set (Fig. 6A right; hold-out split condition). As a control, we also prepared a naive data split condition where the training sample size is the same as in the hold-out split condition but allows overlapping semantic clusters (Fig. 6A left; naive split condition). When we visualized the properties of predicted features in hold-out analysis by transforming them into the previous UMAP embedding space (Fig. 4A), we observed that the predicted features tend to diverge largely from their original clusters and move into other clusters (Fig. 6B).
@@
Fig. 6: Cluster hold-out analysis. (A) Hold-out procedure. For each test sample from a given cluster, all training samples in that cluster are excluded. The naive split (control) uses an equal number of training samples but allows overlap between training and test clusters. (B) Prediction examples. Gray and dark orange points denote training and test samples in the hold-out condition, respectively; light orange points are the predicted latent features, with black lines connecting true features to their predictions. (C) Cluster identification procedure. The predicted features are compared with the cluster centers (average features in each cluster), and the most similar cluster is selected. (D) Cluster identification results. Surrounding panels display CLIP text and vision feature performance in the hold-out and naive split conditions of the four representative semantic clusters. The chance level is $1/40$. (E) Pairwise sample identification procedure. Predicted features are compared with the true test features and with one of the other test features. The sample with features more similar to the prediction is selected. This procedure is repeated for all other test samples, and the proportion of correctly identified true test features is calculated. (F) Pairwise sample identification results. Results are presented similarly to cluster identification, with a chance level of $1/2$.
@@
To quantitatively assess the performance of the feature translator, we employed two identification metrics. The first metric is cluster identification accuracy. Cluster identification accuracy focuses on evaluating the translator’s ability to predict features that correctly identify the semantic cluster to which a test sample belongs (Fig. 6C). In this analysis, we calculated the similarity between the predicted features of test samples and the average features of the training samples within each semantic cluster. The accuracy is then calculated as the percentage of predicted features that successfully identify the original semantic cluster of their corresponding test samples in the latent feature space.
@@
The cluster identification accuracy in the hold-out split condition exhibited a substantial drop compared to the naive split condition across all semantic clusters (Fig. 6D). Notably, the cluster identification accuracy was frequently 00% in the hold-out split condition (see Fig. A6B for all cluster results). These results expose a severe limitation of the translator when dealing with novel semantic clusters absent from the training set. This suggests that the CLIP feature translator primarily functions as a “classifier;” its prediction (translation) heavily relies on predefined semantic features used in a training phase rather than generalizing to new semantic categories.
@@
The second metric is pairwise sample identification accuracy, a commonly used metric in the evaluation of feature prediction and reconstruction performance (Beliy et al., 2019; Shen et al., 2019a, b; Mozafari et al., 2020; Qiao et al., 2020; Ren et al., 2021; Gaziv et al., 2022; Takagi and Nishimoto, 2023a; Ozcelik and VanRullen, 2023; Scotti et al., 2023; Denk et al., 2023; Koide-Majima et al., 2024). This analysis assesses whether the translated features for a given test sample are more similar to its actual features than to those of a randomly chosen sample in the test set (Fig. 6E). The accuracy is calculated as the average winning rate of the predicted features against all the test samples, reflecting how often the predicted features are closer to the correct sample than to any alternative.
Intriguingly, even though the feature translator completely failed to identify the true cluster in the hold-out split condition, pairwise sample identification accuracy often exceeded chance across clusters even in the hold-out split condition (Fig. 6F). This discrepancy arises because pairwise identification is based on relative similarity: if the translated features are “less wrong” for the true sample than for another test sample, they will still be deemed a match. As a result, merely success by a smaller margin than the alternatives can inflate the overall identification score, giving a misleading impression of the translator’s ability to capture new clusters. This observation suggests that relying on pairwise identification accuracy alone, a common practice in many studies, may overestimate reconstruction performance. This issue is further discussed in “Caveat with evaluation by pairwise identification”.
When applying the hold-out split procedure to the full reconstruction pipeline, StableDiffusionReconstruction showed noticeable degradation, whereas Brain-Diffuser remained more robust (Fig. A7). This robustness arises because, even in the hold-out split, the training and test stimuli remain highly similar in the NSD (Fig. A7A). As a result, Brain-Diffuser reconstructions still capture the general visual layout of the original images, with only minor semantic differences (Fig. A7C: 1-cluster hold-out). However, when removing additional clusters such that the similarity between training and test samples approximates that of a completely independent dataset (Fig. A7A: similarity-matched hold-out), the reconstructions, though still realistic, deviate considerably from the original images, both semantically and visually (Fig. A7C: similarity-matched hold-out). This observation suggests that training–test similarity in the NSD played a crucial role in driving text-guided reconstruction performance. When this similarity is reduced to the level of an independent dataset, the ability to reconstruct meaningful details of unseen images diminishes, confirming that the performance of these methods relies on such overlap.
2.1.4 Failed recovery of a stimulus from its latent features